OpenStudy (anonymous):

Use the binomial formula to find the coefficient of the u^7m^3 term in the expansion of (u+2m)^10 .

hero (hero):

Ever heard of Pascal's Triangle? That would be useful here

OpenStudy (anonymous):

That thing has always confused me >.<

hero (hero):

really? It's probably much easier than you think. You just start with 1....then 1 2 1, then 1 3 3 1, then 1 4 6 4 1....and just keep going until you get to the 10th row.

OpenStudy (anonymous):

Hmm.. I kinda see where you're going with that but I'm not sure how I would use that to find the coefficient of u^7m^3 in (u+2m)^10 expanded.. haha sorry.. I'm dumb >.<

hero (hero):

u^10+10u^9m+45u^8m^2+120u^7m^3..... that's how

hero (hero):

I painfully wrote out the Pascal's Triangle up to 10 levels deep...You can thank me later...

OpenStudy (anonymous):

When you expand \[(x+y)^{n}\] using the binomial theorem, the kth term is: \[\left(\begin{matrix}n \\ k-1\end{matrix}\right)x^{n-(k-1)}y^{k-1}\] using this formula, the u^7m^3 term will be: \[\left(\begin{matrix}10 \\ 3\end{matrix}\right)u^7(2m)^3\] \[\left(\begin{matrix}10 \\ 3\end{matrix}\right) = \frac{10!}{3!(10-3)!} = \frac{10*9*8}{1*2*3} = 120\] \[2^3 = 8\] 120*8 = 960 So the coefficient will be 960.

hero (hero):

How come the co-efficient isn't 120?

OpenStudy (anonymous):

because of the 2m. if it was just m it would be 120.

hero (hero):

Yeah, you're right....forgot there are two variables

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%28u%2B2m%29^10

hero (hero):

But it's not 960 either...according to wolf

OpenStudy (anonymous):

Sorry for delayed response but this site is acting up on me still. Thank you guys so very much for your help, I appreciate it! and thanks Hero for writing out Pascal's Triangle!

OpenStudy (anonymous):

for some reason, they put the m first instead of u, i noticed that too lol. make sure you are looking at the u^7m^3 and not the m^7u^3 term.

hero (hero):

You just wrote the same thing twice, lol

OpenStudy (anonymous):

i switched the powers.

hero (hero):

Oh, I see...

hero (hero):

What do I know....I'm just an amateur

OpenStudy (anonymous):

Way better at math than I Hero, that's for sure >.<

hero (hero):

Yeah, but not good enough obviously